I am having trouble trying to figure out how to find the hypothenuse. Please help with this question and explain to me what was done to get the answer.
Find the hypothenuse of a right triangle whose legs measure 5 and 12.

Answer by dawnk925(6)

If I drew a triangle, there would be three sides. The two legs are the shorter sides and the hypotenuse is the side that is opposite of the ninety degree angle.
We can use pythagorean theorem to figure out the side for the missing hypotenuse.
a^2 + b^2 = c^2
5^2+12^2=c^2
25+144= c^2
169=c^2
NOW TAKE THE SQUARE ROOTS OF BOTH SIDES TO FIND OUT C.
13=c
SO the hypotenuse is 13.